Three cubes, all with edges 2 cm., are stacked on top of the other .What is the surface area of the resulting solid?
A)40 B) 56 C) 48 D) 28 E) 96

Answer by edjones(8007)

6*2=12 area of one side of a stack
12*4=48 area of all 4 sides.
4*2=8 area of top and bottom.
48*8=56
.
Ed